OpenStudy (anonymous):

how do you find the maximum value of the function.. f(x)= -x^2 + 6x -5

OpenStudy (precal):

Put it in the graphing calculator. You should be able to get it to calculate the max for you.

OpenStudy (anonymous):

the vertex is \[-\frac{b}{2a}=-\frac{6}{2\times -1}=3\]replace x by 3 to get the y coordinate

OpenStudy (anonymous):

you get \[y=-3^2+6\times 3-5=-9+18-5=9-5=4\] and that is your maximum